Coefficient of x^5 in f(x) = (x+1)(x+2)..(x+6)??

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
Coeff. of x^5 = b
Sum of roots of f(x) = -b/a = -1 -2 - 3 -4 -5 - 6 = -21
Hence, b = 21
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty